# On a fast consistent selection of nested models with possibly unnormalised probability densities

### v2

@@ 4 Simulation Study
In this section, we assess the efficacy of T-GIC1 and T-GIC2 for models with unnormalized PDFs. In Section 4.1, we study the consistency of MGICE for $N\times t$ PDF. In Sections 4.2 and 4.3, we evaluate the consistency (or otherwise) of T-GIC1 and T-GIC2 and the MGICE for two true models, namely the AR model and the polynomial regression model, each defined on $\mathbb{R}$ and driven by $N\times t$ noise/errors. In Section 4.4, we examine a model with a bvariate von Mises PDF on bounded support. Upon replications, the empirical distribution of selected orders by T-GIC’s are obtained, as well as the average MGICE and its standard deviation (SD) for parameters in the true model.
4.1 $N\times t$ PDF
First, we evaluate the efficacy of the MGICE for unnormalized PDFs. Consider an $N\times t(\alpha,k)$ PDF defined as the product of a normal PDF and a Student t-PDF, given by
@@ 4.1 $N\times t$ PDF
for $-\infty<x<\infty$, where $\alpha>0$, $k>0$, $\theta=(\alpha,k)^{T}$ and $C(\theta)$ is the normalizing constant. $N\times t$ is fat-tailed, with the parameters $\alpha$ and $k$ collectively controlling the tail behavior. Specifically, $\alpha$ represents the scale parameter, and $k$ is the power parameter, which is related to the degrees of freedom $\nu$ of a t-PDF via $k=(\nu+1)/2$.
@@
Considering the affine transformation $X=(Y-\mu)/s$ in Baker (2022), the PDF of $Y$ is $\frac{1}{s}f_{N\times t}(\frac{y-\mu}{s}|\alpha,k)$, with 4 parameters $(\mu,s,\alpha,k)$, where $-\infty<\mu<\infty$ and $s>0$. Let $\{y_{i},i=1,2,\cdots,n\}$ represent $n$ observations and denote $f_{y_{i}}(\theta)=\frac{1}{s}f_{N\times t}(\frac{y_{i}-\mu}{s}|\alpha,k)$. Following routine derivations, we have
@@
We conduct 100 replications to evaluate the performance of the MGICE, with sample size $n=1000,3000,\text{and }5000$, where the true parameter values are set to $(\mu^{*},s^{*},\alpha^{*},k^{*})=(0.3,0.5,0.5,1.5)$. Samples are generated using the acceptance-rejection method, by generating random numbers from the $N(0,1/\alpha)$ distribution. For numerical optimization in MGICE, we use the Adaptive Moment Estimation (Adam) algorithm (Kingma and Ba, 2015) to jointly optimize all parameters. Similar to Baker (2022), for all experiments involving the $N\times t$ PDF, we use the sample mean and standard deviation as starting values for $\mu$ and $s$, and regular initial values for $\alpha$ and $k$, e.g. $\alpha=0.25$ or 1, and $k=1$ or 2. Table 4 presents the average MGICE and its SD, showing good overall consistency of MGICE, although there is still room for improvement for the parameters $s^{*},\alpha^{*},$ and $k^{*}$.
Table 4: The average MGICE and its SD for $N\times t$ PDF.
$N$ Parameter MGICE (SD)
@@
4.2 AR Model with $N\times t$ Noise
Now, we consider an AR model with $N\times t$ distributed noise and use T-GIC method to select the model order. The stationary mean-centered AR model of order p with $N\times t$ noise is given by
@@ 4.2 AR Model with $N\times t$ Noise
where $-\infty<a_{1},\cdots,a_{p},c<\infty$, $s>0$, and $\varepsilon_{t}\sim N\times t(\alpha,k),$ identically and independently.
@@
where $W(x_{t},f_{x_{t}})=-||\nabla_{x_{t}}\log(f_{x_{t}})||^{2}-2\Delta_{x_{t}}\log(f_{x_{t}})$.
Discarding the first sum because $W(x_{t},f_{x_{t}}(\theta))$ for $t=1,\cdots L$ are unavailable in the AR(L) model and denoting $n=N-L$, we have
@@
We conduct 100 replications to obtain the frequency distribution of the selected model orders, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $N=1000,3000,$ and $5000$. The true model order $p^{*}$ is set to 3, with true parameters $(a_{1}^{*},a_{2}^{*},a_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified at $(0.50,-0.25,0.10,3.00,0.50,0.50,1.50)$. In each replication, the first 200 data are discarded to ensure stationarity.
Table 5 shows that T-GIC2 tends to underestimate the model order for smaller $N$ but increasingly selects the correct order as $N$ grows. The underfitting for small $N$ is not unexpected due to the small value of $a_{3}^{*}$. In contrast, T-GIC1 overestimates the model order with high probability even for large $N$. Furthermore, Table 6 reports the average MGICE and its SD for the parameters in the true model, suggesting strong consistency, especially for the AR coefficients and the mean-centering parameter $c$, when $N$ is large.
@@
$N$ Method Selected model order $p$
@@
1000 T-GIC1 0 3 22 13 11 1 6 13 14 17
T-GIC2 1 12 52 9 7 0 4 2 6 7
3000 T-GIC1 0 0 28 6 7 9 9 5 16 20
T-GIC2 0 1 81 11 0 4 1 0 0 2
5000 T-GIC1 0 0 29 10 8 7 9 10 7 20
T-GIC2 0 0 88 5 2 1 2 1 0 1
Table 6: The average MGICE and its SD for the AR(3) model with $N\times t$ noise.
$N$ Parameter MGICE (SD)
@@
4.3 Polynomial Regression Model with $N\times t$ Distributed Errors
Next, we consider a polynomial regression model with $N\times t$ distributed errors and use T-GIC to select the model. The polynomial regression model of degree p is given by
@@ 4.3 Polynomial Regression Model with $N\times t$ Distributed Errors
where $-\infty<\beta_{1},\cdots,\beta_{p},c<\infty$, $s>0$, and $\varepsilon\sim N\times t(\alpha,k),$ identically and independently.
@@
We conduct 100 replications to obtain the frequency distribution of the selected model degree $p$, ranging from 1 to 10, using T-GIC1 and T-GIC2, with sample size $n=1000,3000,$ and $5000$. The true model order $p^{*}$ is set to 3, with true parameters $(\beta_{1}^{*},\beta_{2}^{*},\beta_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified as $(-1.5,2.0,5.0,3.0,0.5,0.5,1.5)$.
Table 7 suggests that both T-GIC1 and T-GIC2 consistently select the correct degree in probability as $n$ increases. Moreover, Table 8 reports the average MGICE and its SD for the parameters in the true model, showing good consistency as $n$ increases, although at a slightly slower rate for $\alpha$.
@@
1000 T-GIC1 0 0 94 6 0 0 0 0 0 0
T-GIC2 0 0 97 3 0 0 0 0 0 0
3000 T-GIC1 0 0 96 4 0 0 0 0 0 0
T-GIC2 0 0 96 4 0 0 0 0 0 0
5000 T-GIC1 0 0 97 3 0 0 0 0 0 0
T-GIC2 0 0 97 3 0 0 0 0 0 0
@@ 4.4 A Bivariate Model with a Von Mises PDF
As a final case, we use T-GIC to select the dimension of the parameter space of a model with unnormalized PDF on bounded support. Consider the bivariate von Mises PDF (Singh et al., 2002) of two circular random variables $X_{1},X_{2}$, given by
[…]
@@
for $0\leq x_{1}$, $x_{2}<2\pi$, where $\kappa_{1}$, $\kappa_{2}\geq 0$, $0\leq\mu_{1}$, $\mu_{2}<2\pi$, $-\infty<\lambda<\infty$, $\theta=(\kappa_{1},\kappa_{2},\mu_{1},\mu_{2},\lambda)^{T}$ and $C(\theta)$ is the normalizing constant. The parameter $\lambda$ quantifies the dependency between two circular random variables $X_{1},X_{2}$. $C(\theta)$ is computationally intractable, involving an infinite sum of Bessel functions.
@@
We conduct 100 replications to evaluate the performance of model selection between $m_{1}$ and $m_{2}$, using T-GIC1 and T-GIC2, with sample size $n=300,500,\text{and }1000$. The true parameter dimension is set to 5, with true parameter $\theta^{*}=(2.0,1.0,1.5,2.5,3.0)^{T}$. Samples are generated using the acceptance-rejection method, by generating a random number from the uniform distribution on $[0,2\pi)\times[0,2\pi)$. For numerical optimization in MGICE, we use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm (Nocedal and Wright, 2006) to jointly optimize all parameters, with all parameters initialized to zero.
In all replications, both T-GIC1 and T-GIC2 correctly select the true parameter dimension 5. Table 9 presents the average MGICE and its SD for parameters in the true model $m_{2}$, demonstrating the strong consistency of MGICE especially when $n$ is large.
@@ 5 Real Data
In this section, we apply T-GIC for model selection with real data across three domains. Now, Baker (2022) has argued that the $N\times t$ PDF provides a more realistic framework for modeling real-world data, especially in finance, and serves as an effective tool for assessing robustness and performing sensitivity analyses. In Sections 5.1 and 5.2, we consider fitting autoregression to some finance data, and polynomial regression to some car data, with $N\times t$ noise/errors, using T-GIC1 and T-GIC2 for model selection. We also compare the results with those based on AIC and BIC with Gaussian noise/errors. In Section 5.3, we consider fitting a bivariate model with von Mises PDF to some wind direction data.
@@ 5.1 Finance Data
We analyze the logged returns on the Financial Times Stock Exchange 100 Index (FTSE 100), covering 9013 observations from January 1986 to April 2021. Figure 1 shows its time plot.
Figure 1: Time plot of logged returns on the FTSE 100 from January 1986 to April 2021.
Baker (2022) assumed, for simplicity, independent data and fitted an $N\times t$ distribution. Instead, we consider an AR model to allow for data dependence. For comparison, we use T-GIC1 and T-GIC2 to select an AR model with $N\times t$ noise in equation (4.4), and AIC and BIC to select an AR model with Gaussian noise, given by
@@
where $\varepsilon_{t}\sim N(0,\sigma^{2})$. Tables 10 and 11 summarize the results. Under Gaussian noise, AIC selects a seventh-order model, while BIC selects a sixth-order model. However, under $N\times t$ noise, T-GIC1 selects a sixth-order AR model, while T-GIC2 selects a second-order AR model, both yielding estimates of $\alpha$ and $k$ similar to Baker (2022). Since AR models offer a convenient framework for prediction, Table 12 presents a comparison of out-of-sample performance between the two fitted models, either chosen by AIC vs BIC or by T-GIC1 vs T-GIC2. In each case, we use the first part of the dataset for fitting and the remaining 100 data for the rolling $m$-step-ahead forecast. It is interesting to observe, for T-GIC1 vs T-GIC2, that the fitted AR(6) model performs better in forecasting for the short term than the fitted AR(2) model but worse for the longer term. However, for AIC vs BIC, the fitted AR(6) performs better than the fitted AR(7) uniformly. Overall, Table 12 suggests that the effect of heavy-tailed innovation kicks in for the longer term forecasting.
@@
$p=6$ (T-GIC1) (-0.0212, -0.0813, -0.0158, -0.0265, -0.0230, -0.0331) 0.0006 0.0129 0.0104 2.0799
$p=2$ (T-GIC2) (-0.0256, -0.0877) 0.0007 0.0129 0.0104 2.0807
@@
$m$ $\varepsilon_{t}\sim N\times t(\alpha,k)$ $\varepsilon_{t}\sim N(0,\sigma^{2})$
@@
1 $8.2026\times 10^{-5}$ $8.1961\times 10^{-5}$ 1.0008 $8.1247\times 10^{-5}$ $8.1534\times 10^{-5}$ 0.9965

### v3

@@ 4 Simulation Study
In this section, we assess the efficacy of MIC1 and MIC2 for models with unnormalized PDFs, in comparison with the three bias-corrected criteria GICc, NCIC1, and NCIC2. In Section 4.1, we study the consistency of MGICE for Baker PDF. In Sections 4.2 and 4.3, we evaluate the consistency (or otherwise) of MIC1 and MIC2 and the MGICE for two true models, namely the AR model and the polynomial regression model, each defined on $\mathbb{R}$ and driven by Baker noise/errors. In Section 4.4, we examine a model with a bvariate von Mises PDF on bounded support. Upon replications, the empirical distribution of selected orders by MIC’s is obtained, as well as the average MGICE and its standard deviation (SD) for parameters in the true model.
4.1 Baker PDF
First, we evaluate the efficacy of MGICE for unnormalized PDFs. Consider the Baker PDF (Baker 2022), $N\times t(\alpha,k)$, defined as the product of a normal PDF and a Student t-PDF:
@@ 4.1 Baker PDF
for $\infty<x<\infty$, where $\alpha>0$, $k>0$, $\theta=(\alpha,k)^{T}$ and $C(\theta)$ is the normalizing constant. Baker PDF is fat-tailed, with the parameters $\alpha$ and $k$ collectively controlling the tail behavior. Specifically, $\alpha$ represents the scale parameter, and $k$ is the power parameter, which is related to the degrees of freedom $\nu$ of a t-PDF via $k=(\nu+1)/2$.
@@
Considering the affine transformation $X=(Y-\mu)/s$ in Baker (2022), the PDF of $Y$ is $\frac{1}{s}f_{N\times t}(\frac{y-\mu}{s}|\alpha,k)$, with 4 parameters $(\mu,s,\alpha,k)$, where $\infty<\mu<\infty$ and $s>0$. Let $\{y_{i},i=1,2,\cdots,n\}$ represent $n$ observations and denote $f_{y_{i}}(\theta)=\frac{1}{s}f_{N\times t}(\frac{y_{i}-\mu}{s}|\alpha,k)$. Following routine derivations, we have
@@
We conduct 100 replications to evaluate the performance of the MGICE, with sample size $n=1000,3000,\text{and }5000$, where the true parameter values are set to $(\mu^{*},s^{*},\alpha^{*},k^{*})=(0.3,0.5,0.5,1.5)$. Samples are generated using the acceptance-rejection method, by generating random numbers from the $N(0,1/\alpha)$ distribution. For numerical optimization in MGICE, we use the Adaptive Moment Estimation (Adam) algorithm (Kingma & Ba 2015) to jointly optimize all parameters. Similar to Baker (2022), for all experiments involving the Baker PDF, we use the sample mean and standard deviation as starting values for $\mu$ and $s$, and regular initial values for $\alpha$ and $k$, e.g. $\alpha=0.25$ or 1, and $k=1$ or 2. Table 3 presents the average MGICE and its SD, showing good overall consistency of MGICE, although there is still room for improvement for the parameters $s^{*},\alpha^{*},$ and $k^{*}$.
Table 3: The average MGICE and its SD for Baker PDF.
NN Parameter MGICE (SD)
@@
4.2 AR Model with Baker Noise
Now, we consider an AR model with Baker distributed noise and use MIC method to select the model order. The stationary mean-centered AR model of order p with Baker noise is given by
@@ 4.2 AR Model with Baker Noise
where $\infty<a_{1},\cdots,a_{p},c<\infty$, $s>0$, and $\varepsilon_{t}\sim N\times t(\alpha,k),$ identically and independently.
@@
[…]
Discarding the first sum because $W(x_{t},f_{x_{t}}(\theta))$ for $t=1,\cdots L$ are unavailable in the AR(L) model and denoting n=NLn=N-L, we have
@@
We conduct 100 replications to obtain the frequency distribution of the selected model orders, ranging from 1 to 10, using GICc, MIC1 and MIC2, with sample size $N=1000,3000,$ and $5000$. The true model order $p^{*}$ is set to 3, with true parameters $(a_{1}^{*},a_{2}^{*},a_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified at (0.50,0.25,0.10,3.00,0.50,0.50,1.50)(0.50,-0.25,0.10,3.00,0.50,0.50,1.50). Here, NCIC1 and NCIC2 are not considered in this experiment, as they are developed under the assumption of independent and identically distributed observations, and their applicability to dependent data in stationary AR models remains unclear. In each replication, the first 200 data are discarded to ensure stationarity.
Table 4 shows that MIC2 tends to underestimate the model order for smaller $N$ but increasingly selects the correct order as $N$ grows, outperforming the other criteria overall. The underfitting for small $N$ is not unexpected due to the small value of $a_{3}^{*}$. In contrast, MIC1 overestimates the model order with high probability even for large $N$. Furthermore, Table 5 reports the average MGICE and its SD for the parameters in the true model, suggesting strong consistency, especially for the AR coefficients and the mean-centering parameter $c$, when $N$ is large.
@@
NN Method Selected model order $p$
@@
1000 GICc 0 13 57 18 6 1 3 2 0 0
MIC1 0 3 22 13 11 1 6 13 14 17
MIC2 1 12 52 9 7 0 4 2 6 7
3000 GICc 0 0 71 16 2 4 3 2 1 1
MIC1 0 0 28 6 7 9 9 5 16 20
MIC2 0 1 81 11 0 4 1 0 0 2
5000 GICc 0 0 74 16 3 2 2 2 0 1
MIC1 0 0 29 10 8 7 9 10 7 20
MIC2 0 0 88 5 2 1 2 1 0 1
Table 5: The average MGICE and its SD for the AR(3) model with Baker noise.
NN Parameter MGICE (SD)
@@
4.3 Polynomial Regression Model with Baker Distributed Errors
Next, we consider a polynomial regression model with Baker distributed errors and use MIC to select the model. The polynomial regression model of degree p is given by
@@ 4.3 Polynomial Regression Model with Baker Distributed Errors
where $\infty<\beta_{1},\cdots,\beta_{p},c<\infty$, $s>0$, and $\varepsilon\sim N\times t(\alpha,k),$ identically and independently.
@@
We conduct 100 replications to obtain the frequency distribution of the selected model degree $p$, ranging from 1 to 10, using GICc, NCIC1, NCIC2, MIC1 and MIC2, with sample size $n=300,500,1000,3000,$ and $5000$. The true model order $p^{*}$ is set to 3, with true parameters $(\beta_{1}^{*},\beta_{2}^{*},\beta_{3}^{*},c^{*},s^{*},\alpha^{*},k^{*})$ specified as (1.5,2.0,5.0,3.0,0.5,0.5,1.5)(-1.5,2.0,5.0,3.0,0.5,0.5,1.5). For NCIC1 and NCIC2, noise samples of the same size are generated from the classically estimated polynomial regression model distribution using the least squares method under the Gaussian errors assumption.
Table 6 suggests that both MIC1 and MIC2 consistently select the correct model degree with increasing sample size. Although MIC2 shows slightly lower accuracy than NCIC2 when the sample size is small, its performance remains satisfactory and becomes comparable to that of NCIC2 as $n$ increases, eventually achieving the best performance among the criteria. Moreover, MIC2 exhibits substantially higher computational efficiency than NCIC2. In addition to the theoretical complexity comparison in Section 3.2, we empirically compare the running time ratio of the two methods across different sample sizes, with running time averaged over 10 simulation replications. The experiments were conducted using Python 3.10 on a computer with an Intel Xeon Gold 6133 CPU. The results, reported in Figure 1, indicate that NCIC2 not only requires more running time than MIC2 but also shows a much faster growth in running time as the sample size increases. Table 7 reports the average MGICE and its SD for the parameters in the true model, showing good consistency as $n$ increases, although at a slightly slower rate for $\alpha$.
Figure 1: Empirical running time ratios between NCIC2 and MIC2 for the polynomial regression model across different sample sizes.
@@
300 GICc 0 0 67 23 4 2 3 0 1 0
NCIC1 8 0 68 10 7 2 2 1 1 1
NCIC2 0 0 96 4 0 0 0 0 0 0
MIC1 0 0 77 14 6 1 2 0 0 0
MIC2 0 0 81 11 5 1 2 0 0 0
500 GICc 0 0 85 9 4 1 1 0 0 0
NCIC1 4 1 66 15 8 1 0 1 3 1
NCIC2 0 0 91 9 0 0 0 0 0 0
MIC1 0 0 84 9 5 1 1 0 0 0
MIC2 0 0 87 7 4 1 1 0 0 0
1000 GICc 0 0 93 7 0 0 0 0 0 0
NCIC1 1 0 78 11 3 2 2 0 2 1
NCIC2 0 0 96 4 0 0 0 0 0 0
MIC1 0 0 94 6 0 0 0 0 0 0
MIC2 0 0 97 3 0 0 0 0 0 0
3000 GICc 0 0 97 3 0 0 0 0 0 0
NCIC1 0 0 77 12 5 2 2 2 0 0
NCIC2 0 0 98 2 0 0 0 0 0 0
MIC1 0 0 96 4 0 0 0 0 0 0
MIC2 0 0 96 4 0 0 0 0 0 0
5000 GICc 0 0 97 3 0 0 0 0 0 0
NCIC1 0 0 86 12 0 1 0 1 0 0
NCIC2 0 0 91 9 0 0 0 0 0 0
MIC1 0 0 97 3 0 0 0 0 0 0
MIC2 0 0 97 3 0 0 0 0 0 0
@@ 4.4 A Bivariate Model with a Von Mises PDF
As a final case, we use MIC to select the dimension of the parameter space of a model with unnormalized PDF on bounded support. Consider the bivariate von Mises PDF (Singh et al. 2002) of two circular random variables $X_{1},X_{2}$, given by
$f(x_{1},x_{2}\mid\theta)=C(\theta)\exp\{$ $\kappa_{1}\cos(x_{1}-\mu_{1})+\kappa_{2}\cos(x_{2}-\mu_{2})$
@@
for $0\leq x_{1}$, $x_{2}<2\pi$, where $\kappa_{1}$, $\kappa_{2}\geq 0$, $0\leq\mu_{1}$, $\mu_{2}<2\pi$, $\infty<\lambda<\infty$, $\theta=(\kappa_{1},\kappa_{2},\mu_{1},\mu_{2},\lambda)^{T}$ and $C(\theta)$ is the normalizing constant. The parameter $\lambda$ quantifies the dependency between two circular random variables $X_{1},X_{2}$. $C(\theta)$ is computationally intractable, involving an infinite sum of Bessel functions.
@@
We conduct 100 replications to evaluate the performance of model selection between $m_{1}$ and $m_{2}$, using GICc, NCIC1, NCIC2, MIC1 and MIC2, with sample size $n=300,500,\text{and }1000$. The true parameter dimension is set to 5, with true parameter $\theta^{*}=(2.0,1.0,1.5,2.5,3.0)^{T}$. Samples are generated using the acceptance-rejection method, by generating a random number from the uniform distribution on $0,2\pi)\times[0,2\pi)$. For NCIC1 and NCIC2, noise samples of the same size are generated from the same uniform distribution. For numerical optimization in MGICE, we use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm (Nocedal & Wright 2006) to jointly optimize all parameters, with all parameters initialized to zero.
In all replications, all criteria correctly select the true parameter dimension 5. However, NCIC2 requires substantially more running time than MIC2, over ten times longer when the sample size exceeds 3000, as shown in Figure  2. Table 8 presents the average MGICE and its SD for parameters in the true model $m_{2}$, demonstrating the strong consistency of MGICE especially when $n$ is large.
Figure 2: Empirical running time ratios between NCIC2 and MIC2 for the bivariate model with a Von Mises PDF across different sample sizes.
@@ 5 Real Data
In this section, we apply MIC for model selection with real data across three domains. Now, Baker (2022) has argued that the Baker PDF provides a more realistic framework for modeling real-world data, especially in finance, and serves as an effective tool for assessing robustness and performing sensitivity analyses. In Sections 5.1 and 5.2, we consider fitting autoregression to some finance data, and polynomial regression to some car data, with Baker noise/errors, using MIC1 and MIC2 for model selection. We also compare the results with those based on AIC and BIC with Gaussian noise/errors. In Section 5.3, we consider fitting a bivariate model with von Mises PDF to some wind direction data.
@@ 5.1 Finance Data
We analyze the logged returns of three stock market indices: the Financial Times Stock Exchange 100 Index (FTSE 100) with 9013 observations from January 1986 to April 2021, the Nikkei Stock Average (Nikkei 225) with 5159 observations from July 2003 to August 2024, and the Standard & Poor’s 500 Index (S&P 500) with 3827 observations from January 2010 to March 2025. Their time plots are displayed in Figure 3.
(a) FTSE 100
(b) Nikkei 225
(c) S&P 500
Figure 3: Time plots of logged returns on the FTSE 100, Nikkei 225, and S&P 500.
Baker (2022) assumed, for simplicity, independent data and fitted a Baker distribution. Instead, we consider an AR model to allow for data dependence. For comparison, we use MIC1 and MIC2 to select an AR model with Baker noise in equation (4.4), and AIC and BIC to select an AR model with Gaussian noise, given by
@@
where $\varepsilon_{t}\sim N(0,\sigma^{2})$. Tables 9 and 10 summarize the results. Under Gaussian noise, AIC selects a seventh-order model, while BIC selects a sixth-order model for the FTSE 100. For the Nikkei 225, both AIC and BIC select a first-order model, while for the S&P 500, both select a ninth-order model. However, under Baker noise, MIC1 selects a sixth-order AR model, while MIC2 selects a second-order AR model for the FTSE 100, yielding estimates of $\alpha$ and $k$ similar to Baker (2022). For the Nikkei 225, MIC1 selects a sixth-order AR model, whereas MIC2 favors a first-order AR model. For the S&P 500, both MIC1 and MIC2 select a first-order AR model.
@@
FTSE 100 $p=6$ (MIC1) (-0.0212, -0.0813, -0.0158, -0.0265, -0.0230, -0.0331) 0.0006 0.0129 0.0104 2.0799
$p=2$ (MIC2) (-0.0256, -0.0877) 0.0007 0.0129 0.0104 2.0807
Nikkei 225 $p=6$ (MIC1) (-0.0445, -0.0020, -0.0260, -0.0157, -0.0303, 0.0031) 0.0008 0.0115 0.1030 1.2176
$p=1$ (MIC2) (-0.0479) 0.0011 0.0116 0.1052 1.2292
S&P 500 $p=1$ (MIC1, MIC2) (-0.0168) 0.0007 0.0525 0.0093 1.9206
Since AR models offer a convenient framework for prediction, Table 11 presents a comparison of out-of-sample performance between the fitted models, either chosen by AIC vs BIC or by MIC1 vs MIC2. In each case, we use the first part of the dataset for fitting and the remaining 100 data for the rolling $m$-step-ahead forecast. It is interesting to observe that, for the FTSE 100, the fitted AR(6) model selected by MIC1 performs better in forecasting for the short term than the fitted AR(2) model selected by MIC2 but worse for the longer term. However, for AIC vs BIC, the fitted AR(6) performs better than the fitted AR(7) uniformly. The results suggest that the effect of heavy-tailed innovation kicks in for the longer term forecasting. For the Nikkei 225, the fitted AR(6) model chosen by MIC2 achieves the best forecasting performance. For the S&P 500, the fitted AR(1) model selected by both MIC1 and MIC2, which involves heavy-tailed innovation, uniformly outperforms the fitted AR(9) model selected by AIC and BIC.
@@
Data $m$ $\varepsilon_{t}\sim N\times t(\alpha,k)$ $\varepsilon_{t}\sim N(0,\sigma^{2})$
@@
FTSE 100 1 $8.2026\times 10^{-5}$ $8.1961\times 10^{-5}$ 1.0008 $8.1247\times 10^{-5}$ $8.1534\times 10^{-5}$ 0.9965
@@
$MSE(AR(1))$ $MSE(AR(6))$ $\frac{MSE(AR(1))}{MSE(AR(6))}$ $MSE(AR(1))$ - -
Nikkei 225 1 $4.3790\times 10^{-4}$ $4.3024\times 10^{-4}$ 1.0178 $4.3970\times 10^{-4}$ - -
2 $4348\times 10^{-4} 4.4$$4128\times 10^{-4}$ 1.0050 $4.4373\times 10^{-4}$ - -
3 $4705\times 10^{-4} 4.4$$4556\times 10^{-4}$ 1.0033 $4.4671\times 10^{-4}$ - -
4 $5225\times 10^{-4} 4.4$$4758\times 10^{-4}$ 1.0104 $4.5128\times 10^{-4}$ - -
5 $4.5423\times 10^{-4}$ $4.5263\times 10^{-4}$ 1.0035 $4.5522\times 10^{-4}$ - -
$MSE(AR(1))$ - - $MSE(AR(9))$ - -
S&P 500 1 $8073\times 10^{-5} - - 8.8$$8697\times 10^{-5} - -$
2 $8.9475\times 10^{-5}$ - - $8.9767\times 10^{-5}$ - -
3 $9.0349\times 10^{-5}$ - - $9.1379\times 10^{-5}$ - -
4 $9.1280\times 10^{-5}$ - - $9.2134\times 10^{-5}$ - -
5 $9.2046\times 10^{-5}$ - - $9.4757\times 10^{-5}$ - -